[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu Jul 20, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Thu Jul 20 12:58:40 UTC 2017


Here is today's summary of NHC computer operations:
--- mello/pablo

----------

TAFB:
1. The Qpid ingest process into AWIPS2 crashed yesterday. This caused GFE  
to freeze up for several minutes until the NCP restarted the process. NCO  
examined the cause and could not find any obvious reason. Mello will follow  
up today. Things normal today.

2. ticket submitted to replace black cartridge on the HSU color printer.

HSU:
1. Forecasters reported that the GOES-W data was out both the MERCATOR and  
SAT_PROJ directories from 18-20Z Wednesday. The data flow failed for both  
our local GOES-W and our data flow from NESDIS.

2. Although data flow has been working since then, TSB submitted an NCO  
helpdesk ticket to restart satellite one server and the global imaging  
software.

3. E-PAC ATCF workstation continues to preform slowly, often requiring  
reboots. NCO ticket was opened to investigate.



Other items:
The GFS upgrade did not cause any obvious problems on Wednesday.

TSB updated the contact sheets on both the TAFB and HSU ops floor to  
include contact info for GOES-R issues are reported,
NHC RFCs:
1. An RFC is open for the NET APP upgrade next week. Doug Gear and Mark  
Sardi are the contact points. Hugh and Mike Brennan are notifying our back  
up sites for situation awareness in cause unexpected problems arise.

2. There is an request in for a cosmetic change to the cone graphic. The  
work has actually been done.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170720/45e1474e/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list